Grade 12Physical Chemistry

One hundred miiliters of 0.8 M cuso4 is electrolytes for 30 min by passing current of 5ampere.Calculate the amt of Cuso4 in the solution

Sakchi
9 Years agoGrade 12

1 Answer

Naina Singh
9 Years ago
equivalent mass ofCuSO4 = molar mass/ valency factor = 161/2
weight = EIt/96500 = 161 * 5* 30 * 60 / 2 * 96500 
                            = 7.50 gm of CuSO4